OpenStudy (awy):

Solve cos^2(w)=-8sin(w) for all solutions 0 <= w < 2pi Give your answers accurate to 2 decimal places, as a list separated by commas

OpenStudy (dumbcow):

cos^2 = 1 - sin^2

OpenStudy (awy):

sin^2-8sin-1

OpenStudy (awy):

i used the qudratic formula and got sin(w) = 8+-sqrt(68)/2

OpenStudy (awy):

arcsin both sides and got w=-0.1234187

OpenStudy (dumbcow):

yes then take inverse sin of both answers

OpenStudy (awy):

only 8-sqrt(68)/2 is in the domain right?

OpenStudy (awy):

which is -.1234187 after you take inverse sin

OpenStudy (dumbcow):

oh shoot yeah you are right

OpenStudy (awy):

so im having trouble finding the rest of the solutions

OpenStudy (awy):

the 2nd one i found is 2pi - -0.1234187 = 6.15976

OpenStudy (dumbcow):

ok there are 2 angles which satisfy \[\sin w = \frac{8 -\sqrt{68}}{2}\]